# Activation Function Design Sustains Plasticity in Continual Learning

### v1

@@ C.3 Adaptive, granular slopes are useful—but need guidance to stay in-band
C.4 Saturation Fraction as metric for ’dead units’.
Regarding Saturation Fraction used in Fig. C1 we acknowledge that recent works, such as Sokar et al. (2023) or Liu et al. (2025), highlighted the ’dormant neuron phenomenon’, and ’GraMa’ (Gradient Magnitude Neural Activity Metric) in deep reinforcement learning, proposing a valuable metric based on a neuron’s average absolute activation (or gradient) relative to its layer-neuron connections to identify consistently underutilized units. While this provides a generalized approach to ’dead units’, our investigation into a diverse array of activation functions—each with unique output ranges and saturation characteristics (e.g., ReLU’s unbounded positive output versus Tanh’s strict [-1, 1] bounds, or ELU’s negative saturation plateau)—necessitated a more concrete definition of what constitutes a ’saturated’ or ’effectively dead’ state based on post-activation values. Therefore, while sharing the goal of identifying inactive units, our saturation fraction metrics employ criteria specific to each activation function family (such as proximity to +/-1 for Tanh, near-zero output for ReLU, or low output magnitude relative to batch statistics for functions like Swish or Leaky ReLU). This specific approach allows us to more precisely capture the distinct ways different activation architectures can lead to or avoid states of unresponsiveness under duress, rather than applying a single relative threshold across functions with fundamentally different output scales and properties. See code for the full criteria used for each activation function.
@@ D.1 Strees Protocol
D.4 The width of the dead band predicts shock sensitivity.
As studied in Sec. 4, the Dead-Band Sidth (DBW) critically predicts how severely a network with any activation function will be impacted by saturation-inducing shocks. As previously defined, the DBW score represents the proportion of a typical pre-activation input range where the magnitude of the activation’s first derivative, $\lvert\varphi^{\prime}(x)\rvert$ falls below certain threshold, $\epsilon<10^{-3}$. A higher Dead-Band Width Score thus indicates that a larger portion of common inputs will drive the activation into a state of significantly diminished gradient flow.
Therefore, such hypothesis states that this analytically derived DBW will positively correlate with experimentally observed adverse saturation dynamics. Specifically, activation functions with a larger DBW are expected to exhibit a greater overall saturation impact (higher average Area Under Saturation Curve, Avg. AUSC) and potentially longer average Saturation Fraction (SF) recovery times following a shock. This is because a wider intrinsic dead-band suggests that pre-activation scaling shocks are more likely to push numerous units into these unresponsive regions, and keep them there, diminishing the network’s ability to efficiently desaturate and resume learning.
Figure D1: Correlation of Dead-Band Width Score with Saturation Recovery Metrics (All Gammas Aggregated). Each point represents an activation function configuration. (Left): Average Area Under Saturation Curve (Avg. AUSC) vs. Dead-Band Width Score. A strong positive correlation (Pearson r=0.83, p=0.0016) is observed. (Middle): Average Saturation Fraction (SF) Recovery Time (for successful recoveries, measured by epochs) vs. Dead-Band Width Score. No significant correlation is found (Pearson r=-0.25, p=0.45). (Right): Average SF Non-Recovery Rate (%) vs. Dead-Band Width Score. A strong positive correlation (Pearson r=0.84, p=0.0013) is observed, indicating functions more prone to saturation are more likely to fail SF recovery."
Figure D1 strongly supports this hypothesis when considering Avg. AUSC (left) and Avg. SF Non-Recovery Rate (right). Activations with a higher analytical Dead-Band Width Score do, in fact experience a greater overall saturation impact (higher AUSC), while also being much more likely to fail to recover their saturation fraction after a shock. However, for many activations, if they manage to recover, they do so very quickly (in 1 epoch on average), and this speed does not seem to strongly depend on their Dead-Band Width Score, as observed by looking at Avg. SF Recovery Time (middle).
Overall, these results suggest that the Dead-Band Width Score is a good predictor of the likelihood of getting stuck in a saturated state and the overall severity of that saturation, but not necessarily of how many steps it takes to get out if you manage to get out quickly.
D.5 Performance recovery time on Case Study 2

### v2

@@ C.2 Adaptive, granular slopes are useful—but need guidance to stay in-band
C.3 Saturation Fraction as metric for ’dead units’.
Regarding Saturation Fraction used in Fig. 1 we acknowledge that recent works, such as Sokar et al. (2023) or Liu et al. (2025), highlighted the ’dormant neuron phenomenon’, and ’GraMa’ (Gradient Magnitude Neural Activity Metric) in deep reinforcement learning, proposing a valuable metric based on a neuron’s average absolute activation (or gradient) relative to its layer-neuron connections to identify consistently underutilized units. While this provides a generalized approach to ’dead units’, our investigation into a diverse array of activation functions—each with unique output ranges and saturation characteristics (e.g., ReLU’s unbounded positive output versus Tanh’s strict [-1, 1] bounds, or ELU’s negative saturation plateau)—necessitated a more concrete definition of what constitutes a ’saturated’ or ’effectively dead’ state based on post-activation values. Therefore, while sharing the goal of identifying inactive units, our saturation fraction metrics employ criteria specific to each activation function family (such as proximity to +/-1 for Tanh, near-zero output for ReLU, or low output magnitude relative to batch statistics for functions like Swish or Leaky ReLU). This specific approach allows us to more precisely capture the distinct ways different activation architectures can lead to or avoid states of unresponsiveness under duress, rather than applying a single relative threshold across functions with fundamentally different output scales and properties. See code for the full criteria used for each activation function.
@@ D.1 Strees Protocol
D.4 Performance recovery time on Case Study 2